Gibson Company, which expects to start operations on January 1, 2018, will sell digital cameras in shopping malls. Gibson has budgeted sales as indicated in the following table. The company expects a 15 percent increase in sales per month for February and March. The ratio of cash sales to sales on account will remain stable from January through March.

Required

  1. Complete the sales budget by filling in the missing amounts.

  2. Determine the amount of sales revenue Gibson will report on its first quarter pro forma income statement.

Complete the sales budget by filling in the missing amounts. (Round intermediate calculations and final answers to 2 decimal places.)

Sales January February March
Cash sales $34,000
Sales on account 116,000
Total budgeted sales $150,000

Determine the amount of sales revenue Gibson will report on its first quarter pro forma income statement. (Round intermediate calculations and final answer to 2 decimal places.)

Sales revenue
Answer #1

1.

Complete the missing amount as follows:

1 Sales: March January February $39,100 $44,965 2 Cash sales $34,000 3 Sales on account $133,400 $116,000 $153,410 $150,000 4

________________________________________________________________

Working as follows:

1 Sales: January February March 2 Cash sales S34,000 =B2*115% =C2*115% 3 Sales on account $116,000 =B3*115% =C3*115% 4 Total

________________________________________________________________________

2.

Compute sales revenue:

A Sales revenue in Q-1 $150,000 2 January 3 February $172,500 4 March $198,375 $520,875 5 Total sales revenue